# The equilibrium properties of obvious strategy profiles

### v1

@@ 3 Main results
Consider a large game $G$ with the player space $(I,\mathcal{I},\lambda)$ being the Lebesgue unit interval. All the players have a common action set $A=\{0,1\}$. For each player $i\in I$, her payoff function is given by
$u_{i}(a,\tau)=i+\bigl{(}a-\tau(1)\bigr{)}^{2}.$
@@
Hence the large game $G$ can be viewed as a mapping $G(i)=u_{i}$, for all $i\in I$. Let $\mathbb{Q}$ be the set of all rational numbers on $\mathbb{R}$, and $g\colon I\rightarrow\mathcal{M}(A)$ a randomized strategy profile of the game $G$ defined as follows:
$g(i)=\left\{\begin{array}[]{rcl}\delta_{1}&&\textrm{if}\ \ i\in\mathbb{Q}\cap[0,1]\\ i\delta_{0}+(1-i)\delta_{1}&&\textrm{if}\ \ i\in[0,1]\backslash\mathbb{Q}.\\ \end{array}\right.$
@@
Since all the players in this game have different payoff functions, the strategy profile $g$ is a symmetric strategy profile. Moreover, the societal summary $s(g)$ of $g$ is $\int_{I}g(i)\operatorname{d\!}\lambda(i)=\tfrac{1}{2}\delta_{0}+\tfrac{1}{2}\delta_{1}$. Thus we have $u_{i}(0,s(g))=u_{i}(1,s(g))$ for each player $i\in I$, which implies that $g$ is a Nash equilibrium of $G$. Therefore, $g$ is a symmetric Nash equilibrium and its symmetrized strategy profile $\widetilde{g}$ coincides with $g$ itself. That is, $g=\widetilde{g}=\overline{g}\circ G$. Clearly, $G\colon I\to\mathcal{U}_{A}$ is everywhere continuous but $g\colon I\to\mathcal{M}(A)$ is nowhere continuous, hence $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ is also nowhere continuous.
@@
Then we define a sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to $G$. For each game $G^{n}$, the player space is given by $I^{n}=\{\frac{k}{n}\colon k=1,\ldots,n\}$ endowed with a counting measure $\lambda^{n}$. All the players have a common action set $A=\{0,1\}$. For each player $i\in I^{n}$, her payoff function is given by $u^{n}_{i}(a,\tau)=u_{i}(a,\tau)$, for all $a\in A$ and $\tau\in\mathcal{M}(A).$ Thus we have $G^{n}(i)=u_{i}$ for all $i\in I^{n}$.
For each $n\in\mathbb{Z}_{+}$, we know that $g$ induces an obvious strategy profile $g^{n}$ that is given by $g^{n}(i)=\overline{g}(G^{n}(i))=\overline{g}(G(i))=g(i)$, for all $i\in I^{n}$. Since all the elements of $I^{n}$ are rational numbers, we have $g^{n}(i)=\delta_{1}$ for all $i\in I^{n}$. Thus if $g^{n}$ is viewed as an $\varepsilon_{n}$-Nash equilibrium of $G^{n}$, then $\varepsilon_{n}$ must be no less than $(1-\frac{1}{n})^{2}$, which does not converge to 00.
Despite the counterexample, Theorem 1 remains powerful, as it can be applied to a broad range of large games discussed in the literature. For instance, it is applicable to large games with finite characteristics (i.e., $G(I)$ is a finite subset of $\mathcal{U}_{A}$). Clearly, $\overline{g}$ is almost everywhere continuous when $G(I)$ is a finite set. Consequently, as demonstrated in Corollary 1 below, any sequence of obvious strategy profiles in corresponding finite-player games forms a convergent sequence of approximate symmetric equilibria. Therefore, for players in large finite-player games, adopting the strategies played by their imagined partners in the large game is asymptotically optimal.
Corollary 1.
Let $G\colon(I,\mathcal{I},\lambda)\to\mathcal{U}_{A}$ be a large game such that $G(I)$ is a finite subset of $\mathcal{U}_{A}$, and let $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite games converging to $G$. Suppose that $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ is a sequence of obvious strategy profiles induced by $\overline{g}$. Then, there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
Remark 1.
Theorem 1 can be generalized to games with action correspondences, where a player’s characteristic comprises a feasible action set and a payoff function. In this context, one can similarly define the concepts of auxiliary mapping and obvious strategy profile. Furthermore, it can be shown that the sequence of induced obvious strategy profiles is convergent if the auxiliary mapping is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$.
Remark 2.
In addition to games with finite characteristics, the continuity condition on $\bar{g}$ is also satisfied in many games with infinite characteristics. For example, when $G\colon I\to\mathcal{U}_{A}$ and $\widetilde{g}\colon I\to\mathcal{M}(A)$ are both continuous mappings, [^12] Here $\mathcal{U}_{A}$ is endowed with the sup-norm topology, while $I$ is equipped with the canonical topology induced by the mapping $G\colon I\to\mathcal{U}_{A}$. Additionally, $\mathcal{M}(A)$ is endowed with the weak topology. it can be observed from Figure 1 that $\overline{g}$ is also continuous. For further discussions on continuous Nash equilibria in continuous games, see Kim (1997). [^13] Kim (1997) adopted the argmax topology on the the characteristic space $\mathcal{U}_{A}$, which differs from the sup-norm topology used in this context.
We conclude this section with the second main result, which examines the equilibrium properties of obvious strategy profiles after the resolution of uncertainty. For any randomized strategy profile $g^{n}\colon I^{n}\to\mathcal{M}(A)$ of the finite-player game $G^{n}$, a pure strategy profile $f^{n}\colon I^{n}\to A$ is said to be a realization of $g^{n}$ if $f^{n}(i)\in\operatorname{supp}g^{n}(i)$ for each $i\in I^{n}$. Let $\Theta_{g^{n}}$ represent the set of all possible realizations of $g^{n}$, and let $\mathbb{P}$ be the induced probability measure on $\Theta_{g^{n}}$. [^14] To be precise, for a randomized strategy profile $g^{n}$, there exist a probability space $(\Omega,\Sigma,\mathbb{P})$ and a mapping $x^{n}\colon I^{n}\times\Omega\to A$ such that the distribution induced from $x^{n}_{i}$ is $g^{n}(i)$. Thus, for each $\omega\in\Omega$, $x^{n}(\omega)$ represents a possible realization of $g^{n}$, and $\mathbb{P}$ is the induced probability measure on the space of realizations.
Now, consider a subset of $\Theta_{g^{n}}$, denoted $\Theta_{g^{n}}^{\varepsilon}$, consisting of all realizations of $g^{n}$ that are (pure strategy) $\varepsilon$-Nash equilibria:
[…]
We are now prepared to present the next main result, demonstrating that the realizations of obvious strategy profiles form a convergent sequence of (pure strategy) approximate equilibria with probability approaching one.
Theorem 2.
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$, let $\overline{g}$ be an auxiliary mapping of $G$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then for any $\varepsilon>0$,
$\lim\limits_{n\to\infty}\mathbb{P}(\Theta_{g^{n}}^{\varepsilon})=1.$
Theorem 2 implies that, for players in a large finite game, adopting the obvious strategy profile will be asymptotically optimal not only ex ante but also ex post, with high probability. The ex post equilibrium property in large finite-player games was first studied by Kalai (2004), who examined the approximate ex post stability of equilibria in large Bayesian games. He showed that if the payoff functions satisfy certain equicontinuity conditions, then realizations of an exact equilibrium in a large Bayesian game are approximately stable with high probability. [^15] While Kalai (2004) focused on games with finite action and type spaces, this result was later generalized to games with general action and type spaces in Carmona and Podczeck (2012) and Deb and Kalai (2015). The main difference between Theorem 2 and Kalai (2004) is that Theorem 2 considers a Nash equilibrium of the limiting large game $G$, which may not yield an exact Nash equilibrium in each finite-player game $G^{n}$. Moreover, as demonstrated in Example 1, the continuity assumption on $\overline{g}$ in Theorem 2 is essential.
4 Further results and discussions
This section is organized into three parts. Subsection 4.1 introduces an equilibrium convergence result, demonstrating that any Nash equilibrium of the large game can be approximated by a convergent sequence of approximate symmetric equilibria derived from the corresponding finite-player games. Subsection 4.2 examines obvious strategy profiles in pure strategies, illustrating that, under specific conditions, such strategy profiles exist and form a convergent sequence. Finally, Subsection 4.3 provides a review of relevant literature from prior studies.
4.1 Convergence of approximate symmetric equilibria
Our main results imply that under certain conditions, every equilibrium distribution $\tau$ of the large game induces a sequence of convergent sequence of approximate symmetric equilibria $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ in the corresponding finite-player games. However, whether the sequence of $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to the equilibrium distribution $\tau$ of the large game remains unknown. Generally, given an equilibrium distribution $\tau$ of the large game, does there exist a sequence of symmetric approximate equilibria $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ of finite-player games such that the societal summaries $s(g^{n})$ converge to $\tau$ as $n$ tends to infinity? We provide an affirmative answer to this question in Proposition 1 below.
@@ 4.1 Convergence of approximate symmetric equilibria
Given a sequence of finite games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to $G$, let $\tau$ be an equilibrium action distribution of $G$. Then, there exists a sequence of symmetric strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ such that $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, $\varepsilon_{n}\to 0$, and $s(g^{n})\to\tau$.
Unlike Theorem 1 and Corollary 1, Proposition 1 does not require additional continuity condition or finite-characteristic condition. The proof of Proposition 1 is based on a slight modification of the proof of Theorem 1; therefore, we omit the full detailed proof here but introduce the main idea of our proof instead. Firstly, we consider the case where the continuity condition of $\overline{g}$ holds in $G$. In this situation, we employ the obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ to approximate $\tau$. Indeed, Theorem 1 shows that $\varepsilon_{n}\to 0$, and moreover, the continuity of $\overline{g}$ implies that $s(g^{n})\to\tau$ (see Step 3 of the proof of Theorem 1 in Subsection 5.3). Then for general cases where $\overline{g}$ is not continuous, by using Lusin’s Theorem, [^16] Lusin’s theorem says that for any measurable function $f$, on a metric space $X$ endowed with a finite Borel measure $\lambda$, and for any $\varepsilon>0$, there exists a continuous function $g$ on $X$ such that $\lambda\{x\in X\,|\,g(x)\neq f(x)\}<\varepsilon$. $\overline{g}$ can be approximated by a sequence of continuous functions $\{\overline{g}^{n}\}_{n\in\mathbb{Z}_{+}}$. Following the earlier steps and the proof of Theorem 1, $\overline{g}^{n}\circ G$ can be approximated by a sequence of obvious strategy profiles $\{g^{n,m}\}_{m\in\mathbb{Z}_{+}}$. Finally, employing a diagonal argument, we obtain a sequence of convergent symmetric approximate equilibria that converges to $\tau$.
As a potential enhancement of Proposition 1, a natural question arises: can we strengthen this convergence result by showing that any Nash equilibrium distribution of $G$ can be approximated by a sequence of (exact) Nash equilibria of the corresponding finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. However, such an exact lower hemicontinuous property does not generally hold, and Example 2 below serves as a counterexample. [^17] A notable feature of this example is that the limit Nash equilibrium of the large game does not take weakly dominated strategies.
Example 2.
Consider a large game $G$ with the player space $I$ being the Lebesgue unit interval. Players have a common action set $A=\{a,b\}$, and a common payoff function $u$ that is given by:
$u(a,\tau)=\tau(a)\bigl{(}0.9-\tau(a)\bigr{)},\quad u(b,\tau)=0.$
Clearly, there is no weakly dominated strategy, and $g(i)\equiv b$ is a Nash equilibrium. Similarly, we obtain a sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ by discretizing $I$.
Claim 1.
The Nash equilibrium $g$ of $G$ cannot be approximated by any sequence of Nash equilibria of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$.
The detailed proof of Claim 1 is in Subsection 5.5. Here, we provide a brief intuition. Suppose $s(g)=(0,1)$ is a limit point of a sequence $\{s(g^{n})\}_{n\in\mathbb{Z}_{+}}$ where $g^{n}$ is a Nash equilibrium of $G^{n}$. Then for sufficiently large $n$, $s(g^{n})$ is close to $s(g)$, implying that most players in $g^{n}$ choose action $b$ with high probability. After resolving uncertainty, most players indeed choose action $b$. Therefore, players who choose action $b$ with high probability would deviate to action $a$ to obtain higher payoffs, leading to a contradiction. The main technical challenge arises due to the involvement of randomized strategies, resulting in some realized payoffs being less than 0. To address this, we establish a one-to-one matching between negative and positive payoffs, ensuring that the sum of all realized payoffs is positive.
Proposition 1 examines the convergence of Nash equilibria in large games, which is related to results in the literature. However, earlier studies primarily focused on a different question: whether a convergent sequence of Nash equilibria approaches a Nash equilibrium in the limit game. This inquiry concerns the upper hemicontinuity of equilibria, explored in earlier works such as Green (1984) and Housman (1988). Recent advancements on this topic include Khan et al. (2013), Qiao and Yu (2014), and He et al. (2017). In contrast, Proposition 1 in our paper investigates a lower hemicontinuous property of equilibria in large games.
4.2 Obvious strategy profiles in pure strategies
An obvious strategy profile induced from an auxiliary mapping typically involves randomized strategies. A potential extension is to explore the existence of such a strategy profile in pure strategies. Notably, if the symmetric equilibrium $\widetilde{g}$ is a pure strategy profile, then the auxiliary mapping $\overline{g}$ is also pure, as is the resulting obvious strategy profile $g^{n}=\overline{g}\circ G^{n}$. We will examine the atomless condition introduced by Mas-Colell (1984), which ensures the existence of a pure strategy symmetric equilibrium in large games.
Definition 8 (Atomless condition).
A large game $G\colon I\to\mathcal{U}_{A}$ is said to satisfy the atomess condition if the induce distribution $\lambda G^{-1}$ on $\mathcal{U}_{A}$ is atomless.
Mas-Colell (1984, Theorem 2) demonstrated that a pure strategy symmetric equilibrium exists in a large game whenever it satisfies the atomless condition and has a finite action set. [^18] Khan and Sun (1995a, b) generalized this existence result to large games with countable action sets by establishing a pure symmetrization theorem, under the atomless condition. Building on this result and Theorem 1, we derive the following corollary. It is important to note that, throughout the remainder of this subsection, we consider games with a finite action set.
Corollary 2.
@@ 4.2 Obvious strategy profiles in pure strategies
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$ satisfying the atomless condition, let $\overline{g}$ be an auxiliary mapping of $G$ induced from a pure strategy symmetric equilibrium $\widetilde{g}$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium in pure strategy, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
The continuity condition is essential in Corollary 2, as the atomless condition is satisfied in Example 1; however, the sequence of obvious strategy profiles does not converge. Corollary 2 can also be proved using the main result from Carmona and Podczeck (2022). Specifically, when $\widetilde{g}$ and $\overline{g}$ are pure, the continuity condition on $\overline{g}$, combined with the mapping theorem (Billingsley (1999, Theorem 2.7)) imply that $\lambda^{n}(G^{n},g^{n})^{-1}=\lambda^{n}(G^{n},\overline{g}\circ G^{n})^{-1}$ converges weakly to $\lambda(G,\overline{g}\circ G)^{-1}=\lambda(G,\widetilde{g})^{-1}$. Consequently, their main theorem asserts that $g^{n}=\overline{g}\circ G^{n}$ constitutes an $\varepsilon_{n}$-Nash equilibrium, where $\varepsilon_{n}\to 0$ as $n$ approaches infinity. However, this proof method is not applicable when $\overline{g}$ is not pure, due to the technical challenges associated with the potential realizations of randomized strategy profiles.
Corollary 2 can be extended to encompass exact symmetric equilibria from topological perspective. Recently, Carmona and Podczeck (2020) demonstrated that when the space of large games is endowed with a strong convergence topology [^19] To be precise, the sequence of games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to $G$ in this topology if (i)$\{\lambda(G^{n})^{-1}\}_{n\in\mathbb{Z}_{+}}$ converges weakly to $\lambda G^{-1}$, and (ii) the sequence of supports of distributions also converges. We refer to this as the “strong convergence topology” to distinguish it from the weak convergence topology. and players’ characteristic space is compact, games with a strict equilibrium constitute a generic (i.e., open and dense) subset of the space of large games. Since a strict equilibrium is both pure and symmetric, the main theorem therein implies that for any sequence of finite games converging to a generic large game in the context of strong convergence topology, there exists a symmetric pure strategy equilibrium $\widetilde{g}$ of $G$ that induces a sequence of convergent symmetric equilibria of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. Note that the adoption of strong convergence topology is crucial for this generic result and cannot be replaced by the weak convergence topology employed in the literature and in this paper (see Definition 7).
@@
Another extension of Corollary 2 considers large games with more general action spaces. For instance, in a large game $G$ with a countable action set, we can establish, drawing on the main results from Khan and Sun (1995a, b), that the sequence of obvious strategy profiles forms convergent approximate equilibria when the auxiliary mapping is almost everywhere continuous.

### v2

@@ 4.1 Failure of convergence without the continuity condition
Consider a large game $G$ where the player space is the Lebesgue unit interval $(I,\mathcal{I},\lambda)$, and all players share a common action set $A=\{0,1\}$. Each player $i\in I$ has a payoff function defined by
$u_{i}(a,\tau)=i+\bigl{(}a-\tau(1)\bigr{)}^{2},$
so the large game $G$ can be represented by the mapping $G(i)=u_{i}$ for all $i\in I$.
@@
Let $\mathbb{Q}$ denote the set of rational numbers in $\mathbb{R}$, and define a randomized strategy profile $g\colon I\to\mathcal{M}(A)$ as follows:
$g(i)=\begin{cases}\delta_{1}&\text{if }i\in\mathbb{Q}\cap[0,1],\\ i\delta_{0}+(1-i)\delta_{1}&\text{if }i\in[0,1]\setminus\mathbb{Q}.\end{cases}$
@@
Since all players have distinct payoff functions, $g$ is a symmetric strategy profile. The associated societal summary is $s(g)=\int_{I}g(i)\,d\lambda(i)=\tfrac{1}{2}\delta_{0}+\tfrac{1}{2}\delta_{1}.$ It follows that $u_{i}\bigl{(}0,s(g)\bigr{)}=u_{i}\bigl{(}1,s(g)\bigr{)}$ for every $i\in I$, so $g$ is a Nash equilibrium. Thus, $g$ is a symmetric Nash equilibrium, and its symmetrized profile $\widetilde{g}$ coincides with $g$, i.e., $g=\widetilde{g}=\overline{g}\circ G$.
Observe that the characteristic mapping $G\colon I\to\mathcal{U}_{A}$ is everywhere continuous, while the strategy profile $g\colon I\to\mathcal{M}(A)$ is nowhere continuous. Hence, the auxiliary mapping $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ is also nowhere continuous.
@@
Next, define a sequence of finite-player games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to $G$. For each $n\in\mathbb{Z}_{+}$, let the player set be $I_{n}=\{k/n\mathrel{\mathop{\ordinarycolon}}k=1,\ldots,n\}$, equipped with the uniform measure $\lambda_{n}(i)=1/n$ for all $i\in I_{n}$. Each player has the same action set $A=\{0,1\}$, and for each $i\in I_{n}$, define the payoff function by $u_{i}^{n}(a,\tau)=u_{i}(a,\tau)$ for all $a\in A$ and $\tau\in\mathcal{M}(A)$, so $G_{n}(i)=u_{i}$ for all $i\in I_{n}$.
For each $n$, the direct strategy profile induced by $\overline{g}$ is given by $g^{n}(i)=\overline{g}\bigl{(}G_{n}(i)\bigr{)}=\overline{g}\bigl{(}G(i)\bigr{)}=g(i)$. Since all elements of $I_{n}$ are rational numbers, we have $g^{n}(i)=\delta_{1}$ for every $i\in I_{n}$. However, this sequence of direct strategy profiles does not form a convergent sequence of approximate Nash equilibria. Indeed, for each $i\in I_{n}$, the best deviation from $\delta_{1}$ yields a payoff gain of at least $(1-\tfrac{1}{n})^{2}$. Therefore, if $g^{n}$ is regarded as an $\varepsilon_{n}$-Nash equilibrium of $G_{n}$, it must be that $\varepsilon_{n}\geq(1-\tfrac{1}{n})^{2}$, which does not converge to zero as $n\to\infty$.
The above reasoning shows that the continuity condition in Theorem 1 is indispensable. Moreover, since $\widetilde{g}(i)$ is a pure strategy for each $i\in I_{n}$, each $g^{n}(i)$ is likewise a pure strategy. As a result, the sequence of realizations of direct strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ also fails to form a convergent sequence of approximate Nash equilibria. This confirms that the continuity assumption is equally essential for Theorem 2.
4.2 Direct strategy profiles in pure strategies
A direct strategy profile induced by an auxiliary mapping generally involves randomized strategies. A natural question is whether such a strategy profile can be constructed in pure strategies. A notable special case arises when the symmetric equilibrium $\widetilde{g}$ of the large game is itself a pure strategy profile. In that case, the auxiliary mapping $\overline{g}$ assigns pure strategies to player characteristics, and the resulting direct strategy profile $g^{n}=\overline{g}\circ G_{n}$ is also composed of pure strategies.
To investigate the existence of symmetric pure strategy equilibria in large games, we adopt the atomless condition introduced by Mas-Colell (1984), which guarantees such equilibria exist when the action set is finite.
Definition 8 (Atomless condition).
A large game $G\colon I\to\mathcal{U}_{A}$ is said to satisfy the atomless condition if the induced distribution $\lambda G^{-1}$ on $\mathcal{U}_{A}$ is atomless.
Mas-Colell (1984, Theorem 2) shows that a pure strategy symmetric equilibrium exists in any large game with a finite action set that satisfies the atomless condition. [^14] Khan and Sun (1995a, b) generalized this existence result to large games with countable action sets by establishing a pure symmetrization theorem under the atomless condition. Building on this result and Theorem 1, we obtain the following proposition. Throughout this subsection, we restrict attention to games with finite action sets.
@@ 4.2 Direct strategy profiles in pure strategies
Given a sequence of finite game $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$ satisfying the atomless condition, let $\overline{g}$ be an auxiliary mapping of $G$ induced from a pure strategy symmetric equilibrium $\widetilde{g}$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of direct strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium in pure strategy, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
The continuity condition on $\overline{g}$ remains essential in Proposition 1. Indeed, Example 1 satisfies the atomless condition, yet the induced sequence of direct strategy profiles fails to converge in the absence of continuity.
@@
Proposition 1 can also be extended to large games with more general action spaces. For example, in a large game with a countable action set, one can show—building on the results of Khan and Sun (1995a, b)—that if the auxiliary mapping is almost everywhere continuous, then the sequence of direct strategy profiles still forms a convergent sequence of approximate equilibria.